# algorithmic

### v2

@@ Abstract
The denoising process of diffusion models can be interpreted as an approximate projection of noisy samples onto the data manifold. Moreover, the noise level in these samples approximates their distance to the underlying manifold. Building on this insight, we propose a novel method to enhance sample generation by aligning the estimated noise level with the true distance of noisy samples to the manifold. Specifically, we introduce a noise level correction network, leveraging a pre-trained denoising network, to refine noise level estimates during the denoising process. Additionally, we extend this approach to various image restoration tasks by integrating task-specific constraints, including inpainting, deblurring, super-resolution, colorization, and compressed sensing. Experimental results demonstrate that our method significantly improves sample quality in both unconstrained and constrained generation scenarios. Notably, the proposed noise level correction framework is compatible with existing denoising schedulers (e.g., DDIM), offering additional performance improvements.
@@ 1 Introduction
Previous studies Rick Chang et al. (2017); Permenter & Yuan (2024) have interpreted the denoising process in diffusion models as an approximate projection onto the data manifold, with the noise level $\sigma$ approximating the distance between noisy samples and the data manifold. This perspective views the sampling process an optimization problem, where the goal is to minimize the distance between noisy samples and underlying data manifold using gradient descent. The gradient direction is approximated by the denoiser output with step size determined by the noise level schedule. However, the denoiser requires an estimate of the noise level as input. We claim that accurately estimating the noise level during the denoising process—essentially the distance to the data manifold—is crucial for convergence and accurate sampling.
The expressive capabilities of diffusion models have also made them a compelling choice for image restoration tasks, where generating high-quality, detailed images is essential (Wang et al., 2024). Diffusion models can be used as an image-prior for capturing the underlying structure of image manifold and have shown significant promise for constrained generation such as image restoration (Song et al., 2022; Chung et al., 2023a; Wu et al., 2024). Plug-and-play methods were proposed to utilize pre-trained models without the need for extensive retraining or end-to-end optimization for linear inverse problems such as super-resolution, inpainting and compressed sensing (Chung et al., 2023a; Dou & Song, 2024; Wang et al., 2023). These methods can be interpreted as alternating taking gradient steps towards the constraint set (linear projection for linear inverse problems) and the image manifold (noise direction estimated by the learned denoiser) to find their intersection. However, they may suffer from inconsistency issues if, after each projection step, the noise level no longer approximates distance (Figure 2). Thus correcting the noise level after each step could increase the accuracy of image restoration tasks.
@@
In this work, we propose a novel noise level correction method to refine the estimated noise level and enhance sample generation quality. Our approach introduces a noise level correction network that aligns the estimated noise level of noisy samples more closely with their true distance to the data manifold. By dynamically adjusting the sampling step size based on this corrected noise level estimation, our method improves the sample generation process, significantly enhancing the quality of generated data. Furthermore, our approach integrates seamlessly with existing denoising scheduling methods, such as DDPM (Denoising Diffusion Probabilistic Models) (Ho et al., 2020), DDIM (Denoising Diffusion Implicit Models) (Song et al., 2021a), and EDM (Karras et al., 2022). Furthermore, we extend the application of noise level correction to various image restoration tasks, showing its ability to improve the performance of diffusion-based models such as DDNM (Denoising Diffusion Nullspace Model) in Wang et al. (2023). Our method achieves improved results across tasks including inpainting, deblurring, super-resolution, colorization, and compressed sensing, as illustrated in Figure 1. Additionally, we introduce a parameter-free lookup table as an approximation of the noise level correction network, providing a computationally efficient alternative for improving the performance of unconstrained diffusion models. In summary, our contributions are:
@@
  • Through extensive experiments, we demonstrate that the proposed noise level correction method consistently provides additional performance gains when applied on top of various denoising methods.
@@ 2.1 Diffusion Models
$z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,\text{~{}where~{}}\epsilon\sim\mathcal{N}(0,I),$ (1)
where the noise schedule $\alpha_{t}$ controls the amount of noise injected at each step. Typically, $01>\alpha_{1}>\alpha_{2}>\dots>\alpha_{T}>$ 0, ensuring that $p(z_{T})\sim\mathcal{N}(0,I)$ for large enough $T$ (Ho et al., 2020). For mathematical convenience, a reparameterization is often employed, defining new variables $x_{t}=z_{t}/\sqrt{\alpha_{t}}$, which results in:
@@
Where $\sigma_{t}$ denotes the noise level. Note that the diffusion process is originally presented in variable $z_{t}$, we use the formulation $x_{t}=\frac{z_{t}}{\sqrt{\alpha_{t}}}$ to simplify the forward and reverse denoising processes. A similar formulation can be found in Song et al. (2021b); Karras et al. (2022).
@@
where $\epsilon_{\theta}$ is the predicted noise at step $t$. This iterative process continues until $x_{0}$ is obtained, which represents a denoised sample. On the other hand, the randomized DDPM Ho et al. (2020) follows the update rule:
@@ 2.2 Additional Related Works
Diffusion models have gained significant attention for their ability to learn complex data distributions, excelling in diverse applications such as image generation (Ho et al., 2020; Arechiga et al., 2023), audio synthesis (Kong et al., 2021), and robotics (Chi et al., 2023). Several methods have been proposed to improve the quality of generated samples while reducing the number of iterations. These include techniques such as distillation to reduce sampling steps (Nichol & Dhariwal, 2021), progressive distillation (Salimans & Ho, 2022), consistency models (Song et al., 2023), and improved design space of diffusion models (Karras et al., 2022). On the theoretical side, significant research has explored the non-asymptotic convergence rates of various diffusion samplers, including DDPM (Benton et al., 2024) and DDIM (Li et al., 2024b), contributing to a deeper understanding of the optimization processes underlying diffusion-based models.
@@
Diffusion models have also demonstrated effectiveness in image restoration tasks, including super-resolution (Chung et al., 2023a), inpainting (Kawar et al., 2022), deblurring, and compressed sensing (Wang et al., 2023). These tasks often require strict adherence to data consistency constraints, making diffusion models particularly suitable for addressing such challenges. Prior works such as DDRM (Kawar et al., 2022) and DDNM (Wang et al., 2023) pioneered the use of diffusion models to solve linear inverse problems by projecting noisy samples onto the subspace of solutions that satisfy linear constraints. Other methods have employed alternative approaches, such as computing full projections by enforcing linear constraints or using the gradient of quadratically penalized constraints (Chung et al., 2022; 2023b). Recent work has extended the application of diffusion models to more complex non-convex constraint functions. In these cases, iterative methods such as gradient descent are utilized to guide the sampling process toward satisfying the constraints, as demonstrated in Universal Guidance (Bansal et al., 2023). Furthermore, a provably robust framework for score-based diffusion models applied to image reconstruction was introduced by Xu & Chi (2024), offering robust performance in handling nonlinear inverse problems while ensuring consistency with the observed data.
@@ 3.1 Noise Level as Distance from Noisy Samples to the Manifold
This work builds on the insight that denoising in diffusion models can be interpreted as an approximate projection onto the support of the training-set distribution. Previous studies Rick Chang et al. (2017); Permenter & Yuan (2024) have established this connection. The distance function of a set $\mathcal{K}\subset\mathbb{R}^{n}$, denoted as $\text{dist}_{\mathcal{K}}(x)$, is defined as the minimum distance from a point $x$ to any point $x_{0}\in\mathcal{K}$:
@@
Permenter & Yuan (2024) introduces the following relative-error model, inspired by the manifold hypothesis (Bengio et al., 2013; Fefferman et al., 2016) as well as from the fact that the ideal/optimal denoiser is the gradient of a smoothed distance function:
@@
Leveraging this assumption, we outline the following theorem informally, which summarizes results from Permenter & Yuan (2024) (for formal proofs refer to Permenter & Yuan (2024)):
@@
Theorem 3.1 demonstrates that the estimated clean sample generated by the denoising process $\hat{x}_{0|t}=x_{t}\sigma_{t}\epsilon_{\theta}(x_{t},\sigma_{t}))$ serves as an approximation of the projection of the noisy sample $x_{t}$ onto the manifold $\mathcal{K}$. It also establishes that throughout the sampling process, the distance of a noisy sample $x_{t}$ onto manifold $\mathcal{K}$, can be approximated by the noise level $\sqrt{n}\sigma_{t}$. Given that $\sigma_{t}$ and $\epsilon_{\theta}$ satisfies Assumption 3.1, the DDIM process generates a sequence $d_{t}$ with monotonically decreasing distance to the manifold by gradient descent on the squared-distance function to the manifold. Specifically, the distance of each noisy sample from the manifold is determined by the noise level, $\text{dist}_{\mathcal{K}}(x_{t})=\sqrt{n}\sigma_{t}$, while the gradient $\nabla f(x)$, which is the direction of projection of $x_{t}$ onto the manifold, is estimated by the noise vector $\nabla\text{dist}_{\mathcal{K}}(x_{t})\approx\epsilon^{\ast}_{\theta}(x_{t},\sigma_{t})/\sqrt{n}$.
@@
Figure 2: Denoising approximates projection. (a) Denoising process using DDIM. (b) Constrained denoising process viewed as an alternative projection. Note that in both (a) and (b), the estimated noise level $\sigma_{t}$ does not always match the distance $\text{dist}_{\mathcal{K}}(x_{t})$. (c) Constrained denoising process with noise level correction (NLC). By replacing the prior noise level $\sigma_{t}$ with the more accurate noise level $\hat{\sigma}_{t}$, the projection more closely aligns with the manifold $\mathcal{K}$.
However Theorem 3.1 relies on two key assumptions: (1) a denoiser $\epsilon_{\theta}(x_{t},\sigma_{t})$ satisfying the relative projection-error assumption, and (2) a noise schedule that decrease slowly enough so that for all $t$, we have $\text{dist}_{\mathcal{K}}(x_{t})\approx\sqrt{n}\sigma_{t}$. For sample generation, cumulative errors introduced during the denoising process can lead to biases due to the imperfections of the denoiser Ning et al. (2023). These errors become particularly significant by the final steps, where deviations in the denoising process result in a large mismatch between the true distance $\text{dist}_{\mathcal{K}}(x_{t})$ and the estimated noise level $\sqrt{n}\sigma_{t}$, as illustrated in Figure 2a. As observed, a mismatch between the true distance and the estimated noise level can lead to the clean image estimation using eq. 5 falling outside the manifold $\mathcal{K}$. In constrained sample generation tasks, such as image restoration and guided sample generation, this issue is further exacerbated. As shown in Figure 2b, deviations introduced by guidance terms or projection steps onto the constraint $\mathcal{C}$ amplify the discrepancy between the estimated noise level $\sqrt{n}\sigma_{t}$ and the actual distance $\text{dist}_{\mathcal{K}}(x_{t})$. These deviations are particularly impactful in the later denoising steps and can prevent the reconstructed sample from accurately lying on the data manifold $\mathcal{K}$. Additional details are provided in Appendix A.
@@ 3.2 Noise Level Correction
For $\text{dist}_{\mathcal{K}}(x_{t})$, calculating the ground-truth distance to the manifold is generally not feasible. Instead, we approximate it using the distance between the noisy sample and its clean counterpart in the forward diffusion process. Specifically, given the noisy samples $x_{t}$ generated by the diffusion process from $x_{0}\in\mathcal{K}$ in eq. 2, we estimate the distance as $\text{dist}_{\mathcal{K}}(x_{t})\approx|x_{t}-x_{0}|$. This approximation is reasonable when the projection of $x_{t}$ onto $\mathcal{K}$ satisfies $\text{proj}_{\mathcal{K}}(x_{t})=\text{proj}_{\mathcal{K}}(x_{0}+\sigma_{t}\epsilon)\approx x_{0}$. As illustrated in Permenter & Yuan (2024), this occurs when $\sigma_{t}$ is small (due to the manifold hypothesis, the random noise $\epsilon$ is orthogonal to the manifold $\mathcal{K}$), and when $\sigma_{t}$ is much larger than the diameter of $\mathcal{K}$.
We introduce a neural network to learn $\hat{r}_{t}=r_{\theta}(x_{t},\sigma_{t})$ for noise level correction. To minimize additional computational costs, we design the noise level correction network $r_{\theta}(\cdot)$ to be small and efficient. It leverages the encoder module of the denoiser’s UNet architecture, followed by compact layers that fully utilize the pre-trained denoiser’s capabilities. As shown in Figure 3, the denoiser network uses a UNet structure to estimate the noise vector (denoising direction) based on the noisy image $x_{t}$ and $\sigma_{t}$. Meanwhile, the noise level correction network utilizes the shared encoder, followed by additional neural network blocks, to predict the residual noise level. We train the noise level correction network $r_{\theta}(\cdot)$ alongside a fixed, pre-trained denoiser $\epsilon_{\theta}(\cdot)$, ensuring coordinated improvement in denoising accuracy. In training, to further enhance the noise level correction network, we introduce a scaling factor $\lambda$ to expand the input-output space of $r_{\theta}(x_{t},\sigma_{t})$ within the diffusion process. The objective function for noise level correction is defined as:
$L_{r_{\theta}}:=\mathbf{E}\left[\|\sqrt{n}\sigma_{t}[1+r_{\theta}(\hat{x}_{t},\sigma_{t})]-\sigma_{t}\lambda\|\epsilon_{t}\|\|\right]$ (14)
$\hat{x}_{t}=x_{0}+\sigma_{t}\lambda\epsilon_{t},~{}\epsilon_{t}\sim\mathcal{N}(0,I),~{}\lambda\sim\mathcal{U}(1-\delta,1+\delta)$ (15)
The scaling factor $\lambda$ is sampled from a uniform distribution $\mathcal{U}(1-\delta,1+\delta)$, with $\delta=0.5$ in our experiment, to control the level of variation introduced into the noise level correction.
Figure 3: Neural Network Architecture

### v3

@@ Abstract
The denoising process of diffusion models can be interpreted as an approximate projection of noisy samples onto the data manifold. Moreover, the noise level in these samples approximates their distance to the underlying manifold. Building on this insight, we propose a novel method to enhance sample generation by aligning the estimated noise level with the true distance of noisy samples to the manifold. Specifically, we introduce a noise level correction network, leveraging a pre-trained denoising network, to refine noise level estimates during the denoising process. Additionally, we extend this approach to various image restoration tasks by integrating task-specific constraints, including inpainting, deblurring, super-resolution, colorization, and compressive sensing. Experimental results demonstrate that our method significantly improves sample quality in both unconstrained and constrained generation scenarios. Notably, the proposed noise level correction framework is compatible with existing denoising schedulers (e.g., DDIM), offering additional performance improvements.
@@ 1 Introduction
Previous studies (Rick Chang et al., 2017; Permenter & Yuan, 2024) have interpreted the denoising process in diffusion models as an approximate projection onto the data manifold, with the noise level $\sigma_{t}$ approximating the distance between noisy samples and the data manifold. This perspective views the sampling process as an optimization problem, where the goal is to minimize the distance between noisy samples and the underlying data manifold using gradient descent. The gradient direction is approximated by the denoiser output with step size determined by the noise level schedule. However, the denoiser requires an estimate of the noise level as input. We claim that accurately estimating the noise level during the denoising process—essentially the distance to the data manifold—is crucial for convergence and accurate sampling.
@@
The expressive capabilities of diffusion models have also made them a compelling choice for image restoration tasks, where generating high-quality, detailed images is essential (Wang et al., 2024). Diffusion models can be used as an image-prior for capturing the underlying structure of image manifold and have shown significant promise for constrained generation, such as image restoration (Song et al., 2022; Chung et al., 2023a; Wu et al., 2024). Plug-and-play methods were proposed to utilize pre-trained models without the need for extensive retraining or end-to-end optimization for linear inverse problems such as super-resolution, inpainting, and compressive sensing (Chung et al., 2023a; Dou & Song, 2024; Wang et al., 2023). These methods can be interpreted as alternating taking gradient steps towards the constraint set (linear projection for linear inverse problems) and the image manifold (noise direction estimated by the learned denoiser) to find their intersection. However, they may suffer from inconsistency issues if, after each projection step, the noise level no longer approximates distance (Figure 2). Thus, correcting the noise level after each step could increase the accuracy of image restoration tasks.
In this work, we propose a novel noise level correction method to refine the estimated noise level and enhance sample generation quality. Our approach introduces a noise level correction network that aligns the estimated noise level of noisy samples more closely with their true distance to the data manifold. By dynamically adjusting the sampling step size based on this corrected noise level estimation, our method improves the sample generation process, significantly enhancing the quality of generated data. Furthermore, our approach integrates seamlessly with existing denoising scheduling methods, such as DDPM (Denoising Diffusion Probabilistic Models) (Ho et al., 2020), DDIM (Denoising Diffusion Implicit Models) (Song et al., 2021a), and EDM (Karras et al., 2022). Furthermore, we extend the application of noise level correction to various image restoration tasks, showing its ability to improve the performance of diffusion-based models such as DDNM (Denoising Diffusion Nullspace Model) by (Wang et al., 2023). Our method achieves improved results across tasks including inpainting, deblurring, super-resolution, colorization, and compressive sensing, as illustrated in Figure 1. Additionally, we introduce a parameter-free lookup table as an approximation of the noise level correction network, providing a computationally efficient alternative for improving the performance of unconstrained diffusion models. In summary, our contributions are:
@@
  • Through extensive experiments, we demonstrate that the proposed noise level correction method consistently provides additional performance gains when applied on top of various denoising methods. Our code is available at https://github.com/Walleclipse/Diffusion-NLC/.
@@ 2.1 Diffusion Models
$z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,$ (1)
where $\epsilon\sim\mathcal{N}(0,I)$ and the noise schedule $\alpha_{t}$ controls the amount of noise injected at each step. Typically, $01>\alpha_{1}>\alpha_{2}>\dots>\alpha_{T}>$ 0, ensuring that $p(z_{T})\approx\mathcal{N}(0,I)$ for large enough $T$ (Ho et al., 2020). For mathematical convenience, a reparameterization is often employed, defining new variables $x_{t}=z_{t}/\sqrt{\alpha_{t}}$, which results in:
@@
Where $\sigma_{t}$ denotes the noise level. Note that the diffusion process is originally presented in variable $z_{t}$, we use the formulation $x_{t}=\frac{z_{t}}{\sqrt{\alpha_{t}}}$ to simplify the forward and reverse denoising processes. A similar formulation can be found in (Song et al., 2021b; Karras et al., 2022).
@@
where $\epsilon_{\theta}$ is the predicted noise at step $t$. This iterative process continues until $x_{0}$ is obtained, which represents a denoised sample. On the other hand, the randomized DDPM (Ho et al., 2020) follows the update rule:
@@ 2.2 Additional Related Works
Diffusion models have gained significant attention for their ability to learn complex data distributions, excelling in diverse applications such as image generation (Ho et al., 2020; Arechiga et al., 2023), audio synthesis (Kong et al., 2021), and robotics (Chi et al., 2023). On the theoretical side, significant research has explored the non-asymptotic convergence rates of various diffusion samplers, including DDPM (Benton et al., 2024) and DDIM (Li et al., 2024b), contributing to a deeper understanding of the optimization processes underlying diffusion-based models.
Few-step Sampling. Several methods have been proposed to improve the quality of generated samples while reducing the number of iterations. The first class of these techniques require modifications to the training process, or training of additional models. They include model distillation (Nichol & Dhariwal, 2021), progressive distillation (Salimans & Ho, 2022) or using consistency models (Song et al., 2023; Kim et al., 2023) to predict the endpoint of the sampling process. They can usually produce samples in fewer than 10 steps but are computationally expensive to train and could suffer from training instabilities. The second class of techniques do not require any additional training but instead modifies the sampling process. These include faster samplers known as EDM (Karras et al., 2022), DPM (Lu et al., 2022) or modifications to the sampling noise schedule as in (Sabour et al., 2024; Hang et al., 2024). They do not need additional training and can be directly used on pre-trained models, but this also limits their performance compared to training-based methods.
Adaptive noise schedules. Our proposed method can be seen as a way of introducing additional computation in the sampling process to adaptively change the sampling noise schedule depending on input images. Noise schedules play an important role during sampling, and (Chen, 2023) showed that different schedules are optimal for different data. To tailor the noise schedule to different inputs, (Sahoo et al., 2024) learns to adapt the noise schedule depending on both the pixel location and input. AdaDiff (Tang et al., 2024) dynamically adjusts the inference computation during sampling using a learned uncertainty estimation module.
@@
Diffusion models for inverse problems. Diffusion models have also demonstrated effectiveness in image restoration tasks, including super-resolution (Chung et al., 2023a), inpainting (Kawar et al., 2022), deblurring, and compressive sensing (Wang et al., 2023). These tasks often require strict adherence to data consistency constraints, making diffusion models particularly suitable for addressing such challenges. Prior work such as DDRM (Kawar et al., 2022) and DDNM (Wang et al., 2023) pioneered the use of diffusion models to solve linear inverse problems by projecting noisy samples onto the subspace of solutions that satisfy linear constraints. Other methods have employed alternative approaches, such as computing full projections by enforcing linear constraints or using the gradient of quadratically penalized constraints (Chung et al., 2022; 2023b). Recent work has extended the application of diffusion models to more complex non-convex constraint functions. In these cases, iterative methods such as gradient descent are utilized to guide the sampling process toward satisfying the constraints, as demonstrated in Universal Guidance (Bansal et al., 2023). Furthermore, a provably robust framework for score-based diffusion models applied to image reconstruction was introduced by (Xu & Chi, 2024), offering robust performance in handling nonlinear inverse problems while ensuring consistency with the observed data.
@@ 3.1 Noise Level as Distance from Noisy Samples to the Manifold
This work builds on the insight that denoising in diffusion models can be interpreted as an approximate projection onto the support of the training-set distribution. Previous studies (Rick Chang et al., 2017; Permenter & Yuan, 2024) have established this connection. The distance function of a set $\mathcal{K}\subset\mathbb{R}^{n}$, denoted as $\text{dist}_{\mathcal{K}}(x)$, is defined as the minimum distance from a point $x$ to any point $x_{0}\in\mathcal{K}$:
@@
(Permenter & Yuan, 2024) introduce the following relative-error model, inspired by the manifold hypothesis (Bengio et al., 2013; Fefferman et al., 2016) as well as from the fact that the ideal/optimal denoiser is the gradient of a smoothed distance function:
@@
Leveraging this assumption, we outline the following theorem informally, which summarizes results from (Permenter & Yuan, 2024) (for formal proofs refer to (Permenter & Yuan, 2024)):
@@
Theorem 3.1 demonstrates that the estimated clean sample generated by the denoising process $\hat{x}_{0|t}=x_{t}-\sigma_{t}\epsilon_{\theta}(x_{t},\sigma_{t}))$ serves as an approximation of the projection of the noisy sample $x_{t}$ onto the manifold $\mathcal{K}$. It also establishes that throughout the sampling process, the distance of a noisy sample $x_{t}$ onto manifold $\mathcal{K}$, can be approximated by the noise level $\sqrt{n}\sigma_{t}$. Given that $\sigma_{t}$ and $\epsilon_{\theta}$ satisfies Assumption 3.1, the DDIM process generates a sequence $d_{t}$ with monotonically decreasing distance to the manifold by gradient descent on the squared-distance function to the manifold. Specifically, the distance of each noisy sample from the manifold is determined by the noise level, $\text{dist}_{\mathcal{K}}(x_{t})=\sqrt{n}\sigma_{t}$, while the gradient $\nabla f(x)$, which is the direction of projection of $x_{t}$ onto the manifold, is estimated by the noise vector $\nabla\text{dist}_{\mathcal{K}}(x_{t})\approx\epsilon^{\ast}_{\theta}(x_{t},\sigma_{t})/\sqrt{n}$.
@@
Figure 2: (a) Standard denoising using DDIM. (b) Constrained denoising interpreted as an alternative projection. In both (a) and (b), the estimated noise level $\sigma_{t}$ may not align with the distance $\text{dist}_{\mathcal{K}}(x_{t})$. (c) Constrained denoising with noise level correction (NLC), where replacing $\sigma_{t}$ with a more accurate estimate $\hat{\sigma}_{t}$ yields projections that better align with the manifold $\mathcal{K}$.
Theorem 3.1 exposes two key assumptions that should be satisfied by the denoising process in order to obtain good samples on $\mathcal{K}$: (1) a denoiser $\epsilon_{\theta}(x_{t},\sigma_{t})$ satisfying the relative projection-error assumption, and (2) a noise schedule that decrease slowly enough so that for all $t$, we have $\text{dist}_{\mathcal{K}}(x_{t})\approx\sqrt{n}\sigma_{t}$. For sample generation, cumulative errors introduced during the denoising process can lead to biases due to the imperfections of the denoiser (Ning et al., 2023). These errors become particularly significant by the final steps, where deviations in the denoising process result in a large mismatch between the true distance $\text{dist}_{\mathcal{K}}(x_{t})$ and the estimated noise level $\sqrt{n}\sigma_{t}$, as illustrated in Figure 2a. As observed, a mismatch between the true distance and the estimated noise level can lead to the clean image estimation using eq. 5 falling outside the manifold $\mathcal{K}$. In constrained sample generation tasks, such as image restoration, this issue becomes more pronounced. As illustrated in Figure 2b, deviations introduced by guidance terms or projection steps onto the constraint $\mathcal{C}$ can amplify the mismatch between the estimated noise level $\sqrt{n}\sigma_{t}$ and the actual distance $\text{dist}_{\mathcal{K}}(x_{t})$. This discrepancy is especially significant in the later denoising stages and can hinder the reconstructed sample from accurately aligning with the data manifold $\mathcal{K}$. Additional details are provided in Appendix A.
@@ 3.2 Noise Level Correction
For $\text{dist}_{\mathcal{K}}(x_{t})$, calculating the ground-truth distance to the manifold is generally not feasible. Instead, we approximate it using the distance between the noisy sample and its clean counterpart in the forward diffusion process. Specifically, given the noisy samples $x_{t}$ generated by the diffusion process from $x_{0}\in\mathcal{K}$ in eq. 2, we estimate the distance as $\text{dist}_{\mathcal{K}}(x_{t})\approx|x_{t}-x_{0}|$. This approximation is reasonable when the projection of $x_{t}$ onto $\mathcal{K}$ satisfies $\text{proj}_{\mathcal{K}}(x_{t})=\text{proj}_{\mathcal{K}}(x_{0}+\sigma_{t}\epsilon)\approx x_{0}$. As illustrated in (Permenter & Yuan, 2024), this occurs when $\sigma_{t}$ is small (due to the manifold hypothesis, the random noise $\epsilon$ is orthogonal to the manifold $\mathcal{K}$), and when $\sigma_{t}$ is much larger than the diameter of $\mathcal{K}$.
Figure 3: Architecture of the noise level correction network $r_{\theta}(\cdot)$ and the denoiser $\epsilon_{\theta}(\cdot)$.
We introduce a neural network to learn $\hat{r}_{t}=r_{\theta}(x_{t},\sigma_{t})$ for noise level correction. To minimize computational costs, we design the noise level correction network $r_{\theta}(\cdot)$ to be small and efficient. It leverages the encoder module of the denoiser’s UNet architecture, followed by compact layers that fully utilize the pre-trained denoiser’s capabilities. As shown in Figure 3, the denoiser network uses a UNet structure to estimate the noise vector (denoising direction) based on the noisy image $x_{t}$ and $\sigma_{t}$. Meanwhile, the noise level correction network utilizes the shared encoder, followed by additional neural network blocks, to predict the residual noise level. We train the noise level correction network $r_{\theta}(\cdot)$ alongside a fixed, pre-trained denoiser $\epsilon_{\theta}(\cdot)$, ensuring coordinated improvement in denoising accuracy. In training, to further enhance the noise level correction network, we use a randomly sampled scaling factor $\lambda$ to perturb the true noise level $\sigma_{t}$, expanding the input-output space of $r_{\theta}(x_{t},\sigma_{t})$. The objective function for noise level correction is defined as:
$L_{r_{\theta}}:=\mathbf{E}_{x_{0},t,\epsilon,\lambda}\left[(\sqrt{n}\sigma_{t}[1+r_{\theta}(\hat{x}_{t},\sigma_{t})]-\sigma_{t}\lambda\|\epsilon\|)^{2}\right]$ (14)
$\hat{x}_{t}=x_{0}+\sigma_{t}\lambda\epsilon,~{}\epsilon\sim\mathcal{N}(0,I),~{}\lambda\sim\mathcal{U}(1-\delta,1+\delta)$ (15)
The scaling factor $\lambda$ is sampled from a uniform distribution $\mathcal{U}(1-\delta,1+\delta)$, with $\delta=0.5$ in our experiment, to control the level of variation introduced into the noise level correction. By allowing $\lambda$ to vary, e.g., $\lambda\sim\mathcal{U}(1-\delta,1+\delta)$, the true noise level becomes approximately $\lambda\sigma_{t}\sqrt{n}$ and $r_{\theta}(\hat{x}_{t},\sigma_{t})$ adapts to capture this variation. This approach covers a broader output space compared to the naive setting $\lambda\equiv 1$, leading to better generalization. For further discussion and an ablation study on different $\lambda$ values, please refer to section D.3.